Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (ineedhelpfast12):

http://openstudy.com/study#/updates/55434733e4b0868da1bdb903

OpenStudy (welshfella):

probability of making th e first shot and missing the second = 0.6 * 0.4 = 0.24 probability of both shots = 0.60 * 0.60 = 0.36 expected points = 1*0.24 + 2 * 0.36

OpenStudy (ineedhelpfast12):

wrong question bro

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!